1. Given the set of vertices, determine whether the quadrilateral is a rectangle, rhombus, or square:

** Find the distance and slope or each line segment.

A (−4,2), B (0,3), C (1,−1), D (−3,−2)

This quadrilateral is a… A (Rhombus) B (Rectangle) C (square)
Answer #1

Asswer 1 (-4,2 3- 2 AB) = 177 slope of AB - 04 1 Bel = 117 ; slope of BC = -1 -3 = 48 Tcol=114 slone of CD= 2+1 = IDAI = 117Asswer 1 (-4,2 3- 2 AB) = 177 slope of AB - 04 1 Bel = 117 ; slope of BC = -1 -3 = 48 Tcol=114 slone of CD= 2+1 = IDAI = 117

answered by: ANURANJAN SARSAM